A truck going at a rate of 20 miles per hour takes 6 hours to complete a trip. How many fewer hours would the trip have taken of the truck were traveling at a rate of 30 miles per hour?
Answer by lwsshak3(11628)

A truck going at a rate of 20 miles per hour takes 6 hours to complete a trip. How many fewer hours would the trip have taken of the truck were traveling at a rate of 30 miles per hour?
Travel time=distance/speed
distance=6*20=120 mi
120/30=4 hrs
6-4=2
How many fewer hours would the trip have taken of the truck were traveling at a rate of 30 miles per hour? 2 hrs